# Synergistic eigenanalysis of covariance and Hessian matrices for enhanced binary classification

### v1

@@ Abstract
Covariance and Hessian matrices have been analyzed separately in the literature for classification problems. However, integrating these matrices has the potential to enhance their combined power in improving classification performance. We present a novel approach that combines the eigenanalysis of a covariance matrix evaluated on a training set with a Hessian matrix evaluated on a deep learning model to achieve optimal class separability in binary classification tasks. Our approach is substantiated by formal proofs that establish its capability to maximize between-class mean distance and minimize within-class variances. By projecting data into the combined space of the most relevant eigendirections from both matrices, we achieve optimal class separability as per the linear discriminant analysis (LDA) criteria. Empirical validation across neural and health datasets consistently supports our theoretical framework and demonstrates that our method outperforms established methods. Our method stands out by addressing both LDA criteria, unlike PCA and the Hessian method, which predominantly emphasize one criterion each. This comprehensive approach captures intricate patterns and relationships, enhancing classification performance. Furthermore, through the utilization of both LDA criteria, our method outperforms LDA itself by leveraging higher-dimensional feature spaces, in accordance with Cover’s theorem, which favors linear separability in higher dimensions. Our method also surpasses kernel-based methods and manifold learning techniques in performance. Additionally, our approach sheds light on complex DNN decision-making, rendering them comprehensible within a 2D space.
@@ 1 Introduction
Traditional approaches often rely on analyzing either the covariance matrix (Nagai2020TheCM, ; Minh2017CovariancesIC, ; Serra2014CovarianceOC, ; Lenc2016LearningCF, ; Hoff2011ACR, ; Kuo2002ACE, ; Lam2019HighdimensionalCM, ) or the Hessian matrix (Dawid2021HessianbasedTF, ; Fu2020HesGCNHG, ; Yao2019PyHessianNN, ; Krishnasamy2016HessianSE, ; Wiesler2013InvestigationsOH, ; Byrd2011OnTU, ; Martens2010DeepLV, ) separately to optimize machine learning models. In the field of Evolution Strategies (ESs), a recent work (shir2020covariance, ) explored the relationship between the covariance matrix and the landscape Hessian, highlighting the statistical learning capabilities of ESs using isotropic Gaussian mutations and rank-based selection. While this study provides valuable insights into ESs’ learning behavior, it does not investigate the practical integration of the two matrices.
@@
To tackle this challenge, we present a novel approach that combines the eigenanalysis of the covariance matrix evaluated on a training set with the Hessian matrix evaluated on a deep learning model. By integrating these matrices, our method aims to optimize class separability by simultaneously maximizing between-class mean distance and minimizing within-class variances, which are the fundamental criteria of linear discriminant analysis (LDA) (fisher1936use, ; xanthopoulos2013linear, ). By utilizing these criteria, we leverage the foundational principles that have been established and proven over decades of LDA’s application in various fields like medicine (sharma2012between, ; sharma2008cancer, ; moghaddam2006generalized, ; dudoit2002comparison, ; chan1995computer, ), agriculture (tharwat2017one, ; gaber2015plant, ; rezzi2005classification, ; heberger2003principal, ; chen1998rapid, ), and biometrics (paliwal2012improved, ; yuan2007ear, ; park2005fingerprint, ; wang2004random, ; yu2001direct, ; chen2000new, ; haeb1992linear, ). LDA has been demonstrating the reliability and usefulness of the two criteria as indicators of discriminative power. Thus, by building upon this well-established foundation, our approach will inherit the strengths and reliability that have been demonstrated by LDA. This integrated approach holds the promise of enhancing classification performance beyond conventional methods that treat the two matrices separately.
@@ 2.1 Approach: combining eigenanalysis of covariance and Hessian matrices
    $\text{{Cov}}(\theta_{1},\theta_{2})=\frac{1}{n-1}\sum_{i=1}^{n}(\theta_{1i}-\bar{\theta}_{1})(\theta_{2i}-\bar{\theta}_{2})$
@@
    $\text{BCELoss}=-\sum_{i=1}^{n}\log p_{\theta}(c_{i}\mid x_{i}),$
@@
    $H_{\theta}$ $=\nabla_{\theta}^{2}\text{BCELoss}$
@@
    Unlike LDA, which aims to optimize both criteria simultaneously along a single direction for binary classification, constrained by the limitation that the number of linear discriminants is at most $c-1$ where $c$ is the number of class labels ye2004two , our approach provides more flexibility and control. By working on two separate directions, we specifically focus on minimizing the within-class variances in one direction while maximizing the between-class mean distance in the other direction.
@@
In this subsection, we establish two theorems along with their respective proof sketches that form the theoretical basis for our approach. Full proofs for both theorems are available in Appendix A.
@@
Consider two sets of 1D data points representing two classes, denoted as $C_{1}$ and $C_{2}$, each consisting of $n$ samples. The data in $C_{1}$ and $C_{2}$ follow the same underlying distribution centered around their respective means, $m_{1}$ and $m_{2}$. Here, $C_{2}$ can be understood as a reflection of $C_{1}$ with the axis of reflection positioned at the overall mean, denoted as $m$. Furthermore, the variances of $C_{1}$ and $C_{2}$, are equal, denoted as $s_{1}^{2}=s_{2}^{2}=s_{w}^{2}$, and the combined data from $C_{1}$ and $C_{2}$ has a variance of $s^{2}$. The between-class mean distance, denoted as $d$, represents the separation between the means of $C_{1}$ and $C_{2}$. We establish the following relationship:
$s^{2}=\frac{d^{2}}{4(1-\lambda)}$ (3)
where $\lambda=\frac{s_{w}^{2}}{s^{2}}$ signifies a constant between 0 and 1 that reflects the distribution of the original data when $C_{1}$ and $C_{2}$ are considered as projected representations.
@@
    Simplify the expression for the combined data variance $s^{2}$ to $s_{w}^{2}+\frac{1}{4}d^{2}$.
@@
    Apply the Variance Ratio Preservation Theorem for Projection onto a Vector to relate $s_{w}^{2}$ and $s^{2}$ through $\lambda$.
@@
    Simplify the expression to obtain $s^{2}=\frac{d^{2}}{4(1-\lambda)}$.
@@
    Conclude that maximizing $s^{2}$ maximizes $d^{2}$, fulfilling the theorem’s objective.
@@ 2.2.2 Theorem 2: Maximizing Hessian for Minimizing Within-Class Variance
$\mathrm{H}_{\theta}\approx\frac{1}{\sigma_{\text{post}}^{2}}.$ (4)
@@
    Approximate the Hessian $\mathrm{H}_{\theta}$ as the Fisher information using the expectation of the squared gradient of the log-likelihood (barshan2020relatif, ).
@@
    Derive $\mathrm{H}_{\theta}\approx\frac{1}{\sigma_{\text{post}}^{2}}$.
@@
Theorem 1 and Theorem 2 respectively suggest that maximizing the variance of projected data and the Hessian effectively maximize squared between-class mean distance and minimize within-class variances. These theorems provide the theoretical foundation for the eigenanalysis of the covariance and Hessian matrices, crucial steps in our proposed method for improving class separability based on the two LDA criteria.
@@ 4.1 Assessed dimensionality reduction techniques
In our evaluation, we compare nine distinct dimensionality reduction and data projection techniques, each offering unique insights: principal component analysis (PCA), kernel PCA (KPCA) (scholkopf1997kernel, ), Hessian, uniform manifold approximation and projection (UMAP) (mcinnes2018umap, ), locally linear embedding (LLE) (roweis2000nonlinear, ), linear optimal low-rank projection (LOL) (vogelstein2021supervised, ), linear discriminant analysis (LDA), kernel discriminant analysis (KDA) (mika1999fisher, ), and the proposed approach. PCA involves projection onto the primary two covariance eigenvectors, i.e., it applies KPCA with a linear kernel. Similarly, LDA employs KDA with a linear kernel. The Hessian method projects data onto the leading two Hessian eigenvectors. Notably, for both KPCA and KDA, the best kernels (linear, polynomial, RBF, sigmoid, or cosine similarity) and their associated parameters (kernel coefficient or degree) are determined using grid search tailored to each dataset. The proposed method combines the most relevant eigenvectors derived from both the covariance and Hessian matrices, as specified by Eqs(1, 2).
@@ 4.3 Datasets
Widely recognized benchmark datasets. We evaluate our approach using three widely recognized benchmark datasets for binary classification: the Wisconsin breast cancer dataset (street1993nuclear, ), the heart disease dataset (detrano1989international, ), and the Pima Indians diabetes dataset (smith1988using, ). Prior to applying various dimensionality reduction methods, we enact standard data preprocessing techniques on the original datasets, including handling of missing data, one-hot encoding for categorical variables, and data normalization.
Neural spike train dataset. A spike train is a sequence of action potentials (spikes) emitted by a neuron over time. The neural spike train dataset used in this research consists of recordings from rat’s neurons during drug application from a multi-electrode array (tsai2015high, ; heuschkel2002three, ). The data, comprising 221 records, represents the final dataset after all preprocessing steps suggested by (lazarevich2023spikebench, ). Each record contains 15 time-series features extracted using the ‘tsfresh‘ package (christ2018time, ) and 1 class attribute indicating whether the neuron is non-responsive (0) or responsive (1). Exploratory data analysis revealed an imbalance in the dataset, with 190 non-responsive neurons (86%) and 31 responsive neurons (14%).
@@ 4.4 Results
The results depicted in Figure 1, which pertain to the WBCD dataset, along with the corresponding findings presented in Appendix B for the other three datasets, collectively reaffirm the consistency and robustness of our proposed approach across diverse datasets. The heatmaps shown in Figure 1(b) and Appendix B consistently demonstrate the reduction in squared between-class mean distance as more covariance eigenvectors are incorporated, aligning with our established theoretical framework (Eq(3)). Furthermore, Figure 1(c) and its counterparts in the appendix reveal the ascending order of within-class variances, in sync with the descending order of the Hessian, supporting our theoretical foundation (Eq(4)). Furthermore, Figure 1(d) and its counterparts illustrate the LDA ratio, emphasizing the equal contributions of covariance and Hessian eigenanalyses to class separability. Importantly, the results also imply that the highest LDA ratio is observed for the combination of the first Hessian eigenvector with the first covariance eigenvector. This observation underscores the significance of projecting data onto the combined space of these primary eigenvectors (as outlined in Eqs (1, 2)), forming the core of our proposed method. These consistent empirical results across multiple datasets not only validate our theoretical premises but also endorse the effectiveness of our proposed method in optimizing class separability.
The evaluation results in Figure 2 compare the performance of different data projection methods, including PCA, KPCA, Hessian, UMAP, LLE, LOL, LDA, KDA, and the proposed method, using 5- or 10-fold cross-validation. It is essential to note that we introduced nonlinearity to the comparative analysis by utilizing distinct kernels determined through grid search in KPCA and KDA for each dataset. Notably, our proposed method consistently outperforms all others, securing the highest scores across all datasets and evaluation metrics. This consistent superiority of our approach implies its potential as a valuable tool for improving classification performance in various domains, further highlighting its promise as a robust and effective method for dimensionality reduction and data projection.

### v2

@@ Abstract
Covariance and Hessian matrices have been analyzed separately in the literature for classification problems. However, integrating these matrices has the potential to enhance their combined power in improving classification performance. We present a novel approach that combines the eigenanalysis of a covariance matrix evaluated on a training set with a Hessian matrix evaluated on a deep learning model to achieve optimal class separability in binary classification tasks. Our approach is substantiated by formal proofs that establish its capability to maximize between-class mean distance and minimize within-class variances, particularly under ideal data conditions such as isotropy around class means and dominant leading eigenvalues. By projecting data into the combined space of the most relevant eigendirections from both matrices, we achieve optimal class separability as per the linear discriminant analysis (LDA) criteria. Empirical validation across neural and health datasets consistently supports our theoretical framework and demonstrates that our method outperforms established methods. Our method stands out by addressing both LDA criteria, unlike PCA and the Hessian method, which predominantly emphasize one criterion each. This comprehensive approach captures intricate patterns and relationships, enhancing classification performance. Furthermore, through the utilization of both LDA criteria, our method outperforms LDA itself by leveraging higher-dimensional feature spaces, in accordance with Cover’s theorem, which favors linear separability in higher dimensions. Our method also surpasses kernel-based methods and manifold learning techniques in performance. Additionally, our approach sheds light on complex DNN decision-making, rendering them comprehensible within a 2D space.
@@ 1 Introduction
Traditional approaches often rely on analyzing either the covariance matrix [1, 2, 3, 4, 5, 6, 7] or the Hessian matrix [8, 9, 10, 11, 12, 13, 14] separately to optimize machine learning models. In the field of Evolution Strategies (ESs), a recent work [15] explored the relationship between the covariance matrix and the landscape Hessian, highlighting the statistical learning capabilities of ESs using isotropic Gaussian mutations and rank-based selection. While this study provides valuable insights into ESs’ learning behavior, it does not investigate the practical integration of the two matrices.
@@
To tackle this challenge, we present a novel approach that combines the eigenanalysis of the covariance matrix evaluated on a training set with the Hessian matrix evaluated on a deep learning model. By integrating these matrices, our method aims to optimize class separability by simultaneously maximizing between-class mean distance and minimizing within-class variances, which are the fundamental criteria of linear discriminant analysis (LDA) [16, 17]. By utilizing these criteria, we leverage the foundational principles that have been established and proven over decades of LDA’s application in various fields like medicine [18, 19, 20, 21, 22], agriculture [23, 24, 25, 26, 27], and biometrics [28, 29, 30, 31, 32, 33, 34]. LDA has been demonstrating the reliability and usefulness of the two criteria as indicators of discriminative power. Thus, by building upon this well-established foundation, our approach will inherit the strengths and reliability that have been demonstrated by LDA. This integrated approach holds the promise of enhancing classification performance beyond conventional methods that treat the two matrices separately.
Clustering algorithms [35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46] also adhere to principles similar to those of LDA criteria. Just as our proposed method leverages the LDA criteria for optimizing class separability by maximizing between-class mean distance and minimizing within-class variances, clustering techniques aim to enhance cluster separability by maximizing both inter-cluster dissimilarity (the concept of separation) and intra-cluster similarity (the concept of compactness) [47, 48, 49, 50]. This alignment highlights the effectiveness and utility of LDA criteria in various data analysis techniques, demonstrating that these principles are broadly applicable and beneficial for achieving distinct and compact data groupings.
@@ 2.1 Approach: combining eigenanalysis of covariance and Hessian matrices
    $\text{{Cov}}(\theta_{1},\theta_{2})=\frac{1}{n-1}\sum_{i=1}^{n}(\theta_{1i}-\bar{\theta}_{1})(\theta_{2i}-\bar{\theta}_{2})$ (1)
@@
    $\text{BCELoss}=-\sum_{i=1}^{n}\log p_{\theta}(c_{i}\mid x_{i}),$ (2)
@@
    $H_{\theta}$ $=\nabla_{\theta}^{2}\text{BCELoss}$ (3)
@@
    Unlike LDA, which aims to optimize both criteria simultaneously along a single direction for binary classification, constrained by the limitation that the number of linear discriminants is at most $c-1$ where $c$ is the number of class labels [51], our approach provides more flexibility and control. By working on two separate directions, we specifically focus on minimizing the within-class variances in one direction while maximizing the between-class mean distance in the other direction.
@@
In this subsection, we establish two theorems along with their respective proof sketches that form the theoretical basis for our approach. Full proofs for both theorems are available in A.
@@
Consider two sets of 1D data points representing two classes, denoted as $C_{1}$ and $C_{2}$, each consisting of $n$ samples. The data in $C_{1}$ and $C_{2}$ follow the same underlying distribution centered around their respective means, $m_{1}$ and $m_{2}$. Here, $C_{2}$ can be understood as a reflection of $C_{1}$ with the axis of reflection positioned at the overall mean, denoted as $m$. Furthermore, the variances of $C_{1}$ and $C_{2}$ are equal, denoted as $\sigma_{1}^{2}=\sigma_{2}^{2}=\sigma_{w}^{2}$, and the combined data from $C_{1}$ and $C_{2}$ has a variance of $\sigma^{2}$. The between-class mean distance, denoted as $d$, represents the separation between the means of $C_{1}$ and $C_{2}$. We establish the following relationship:
$\sigma^{2}=\frac{d^{2}}{4(1-\lambda)}$ (6)
where $\lambda=\frac{\sigma_{w}^{2}}{\sigma^{2}}$ signifies a constant between 0 and 1 that reflects the distribution of the original data when $C_{1}$ and $C_{2}$ are considered as projected representations.
@@
    Simplify the expression for the combined data variance $\sigma^{2}$ to $\sigma_{w}^{2}+\frac{1}{4}d^{2}$.
@@
    Apply the Variance Ratio Preservation Theorem for Projection onto a Vector to relate $\sigma_{w}^{2}$ and $\sigma^{2}$ through $\lambda$.
@@
    Simplify the expression to obtain $\sigma^{2}=\frac{d^{2}}{4(1-\lambda)}$.
@@
    Conclude that maximizing $\sigma^{2}$ maximizes $d^{2}$, fulfilling the theorem’s objective.
In the context of multidimensional data, extending Theorem 1 involves considering data points as $D$-dimensional vectors, where the dataset consists of two subsets, $C_{1}=\{\vec{x}_{i}\in\mathbf{R}^{D}\mid i=1,2,\ldots,n\}$ with covariance $Cov_{1}$, and $C_{2}=\{\vec{y}_{j}\in\mathbf{R}^{D}\mid j=1,2,\ldots,n\}$ with covariance $Cov_{2}$. The theorem, however, is supposed to hold for the projection of the data onto an arbitrary 1D subspace of the full D-dimensional data space, and to imply that maximizing $\sigma^{2}$ by choosing the direction of this subspace appropriately will also be the direction that maximizes the difference between the projected sample means. A general case in which it holds is when the individual distributions are isotropic about their means, meaning the covariances $Cov_{1}$ and $Cov_{2}$ are proportional to the identity matrix in $\mathrm{D}$ dimensions as mathematically shown in B.
@@ 2.2.2 Theorem 2: Maximizing Hessian for minimizing within-class variance
$\mathrm{H}_{\theta}=\frac{1}{\sigma_{\text{post}}^{2}}.$ (7)
@@
    Approximate the Hessian $\mathrm{H}_{\theta}$ as the Fisher information using the expectation of the squared gradient of the log-likelihood [52].
@@
    Derive $\mathrm{H}_{\theta}=\frac{1}{\sigma_{\text{post}}^{2}}$.
@@
Theorem 1 and Theorem 2 respectively suggest that maximizing the variance of projected data and the Hessian effectively maximize squared between-class mean distance and minimize within-class variances, contingent upon the ideal data conditions discussed in Section 3. These theorems provide the theoretical foundation for the eigenanalysis of the covariance and Hessian matrices, crucial steps in our proposed method for improving class separability based on the two LDA criteria.
3 Ideal dataset conditions for the optimal performance of the proposed method
To ensure the optimal performance of the proposed method, it is crucial to consider certain ideal conditions for the dataset. These conditions represent ideal scenarios that can help achieve optimal performance and support the theoretical foundations of the method. The key conditions are as follows:
  1. 1.
    Isotropic distributions around class means:
    The theoretical underpinnings of the proposed method, particularly in the application of Theorem 1 to multidimensional data, benefit from the dataset subsets corresponding to each class being approximately isotropic around their respective means, as mathematically shown in B. While perfect isotropy—where the covariance matrices of the subsets are proportional to the identity matrix in $D$ dimensions—is rare in practice, achieving near-isotropic conditions helps ensure the validity of the theorem to higher dimensions and the method’s ability to leverage the relationship between covariance and class separability.
    In more practical terms, datasets tend to support isotropism under the following conditions:
    1. (a)
      Normalized data: Normalizing the data, typically through techniques such as z-score normalization, ensures that each attribute has a mean of zero and a standard deviation of one. This process equalizes variance in all directions, which approximates the isotropic condition.
    2. (b)
      Low correlation between attributes: Datasets with low correlation between attributes produce covariance matrices with minimized off-diagonal elements. This condition approximates isotropy, which is characterized by covariance matrices being proportional to the identity matrix.
  2. 2.
    Dominance of the first eigenvalues in covariance and Hessian matrices:
    The effectiveness of the proposed method is heavily reliant on the information provided by the leading eigenvectors derived from the covariance and Hessian matrices. For these matrices to significantly contribute to the method’s success, their first eigenvalues need to be sufficiently dominant. A dominant first eigenvalue indicates that the corresponding eigenvector captures the most significant variance or curvature in the data, making it a crucial direction for class separation.
    The presence of a dominat first eigenvalue can be assessed from the eigenspectra of the two matrices, which display the eigenvalues on a log scale. A dominant first eigenvalue is characterized by the eigenvalues being spread over many orders of magnitude, often with approximately uniform spacing between them. When a dominant first eigenvalue is present in the covariance matrix, it suggests a dominant underlying structure or trend in the data. When this occurs in the Hessian matrix, it characterizes a ”sloppy” model [53, 54, 55, 56, 57, 58, 59], which has a few stiff directions (significant eigenvalues) and many sloppy directions (small eigenvalues).
@@ 5.1 Assessed dimensionality reduction techniques
In our evaluation, we compare nine distinct dimensionality reduction and data projection techniques, each offering unique insights: principal component analysis (PCA), kernel PCA (KPCA) [60], Hessian, uniform manifold approximation and projection (UMAP) [61], locally linear embedding (LLE) [62], linear optimal low-rank projection (LOL) [63], linear discriminant analysis (LDA), kernel discriminant analysis (KDA) [64], and the proposed approach. PCA involves projection onto the primary two covariance eigenvectors, i.e., it applies KPCA with a linear kernel. Similarly, LDA employs KDA with a linear kernel. The Hessian method projects data onto the leading two Hessian eigenvectors. Notably, for both KPCA and KDA, the best kernels (linear, polynomial, RBF, sigmoid, or cosine similarity) and their associated parameters (kernel coefficient or degree) are determined using grid search tailored to each dataset. The proposed method combines the most relevant eigenvectors derived from both the covariance and Hessian matrices, as specified by Eqs(4, 5).
@@ 5.3 Datasets
Widely recognized benchmark datasets. We evaluate our approach using three widely recognized benchmark datasets for binary classification: the Wisconsin breast cancer dataset [65], the heart disease dataset [66], and the Pima Indians diabetes dataset [67]. Prior to applying various dimensionality reduction methods, we enact standard data preprocessing techniques on the original datasets, including handling of missing data, one-hot encoding for categorical variables, and data normalization. Specifically, we apply z-score normalization to ensure the data has zero mean and unit variance, supporting the requirement for the data to be isotropic as suggested in Section 3.
Neural spike train dataset. A spike train is a sequence of action potentials (spikes) emitted by a neuron over time. The neural spike train dataset used in this research consists of recordings from rat’s neurons during drug application from a multi-electrode array [68, 69]. The data, comprising 221 records, represents the final dataset after all preprocessing steps suggested by [70]. Each record contains 15 time-series features extracted using the ‘tsfresh‘ package [71] and 1 class attribute indicating whether the neuron is non-responsive (0) or responsive (1). Exploratory data analysis revealed an imbalance in the dataset, with 190 non-responsive neurons (86%) and 31 responsive neurons (14%).
@@ 5.4 Results
The results depicted in Figure 1, which pertain to the WBCD dataset, along with the corresponding findings presented in C for the other three datasets, collectively reaffirm the consistency and robustness of our proposed approach across diverse datasets. The heatmaps shown in Figure 1(b) and C consistently demonstrate the reduction in squared between-class mean distance as more covariance eigenvectors are incorporated, aligning with our established theoretical framework (Eq(6)). Furthermore, Figure 1(c) and its counterparts in the appendix reveal the ascending order of within-class variances, in sync with the descending order of the Hessian, supporting our theoretical foundation (Eq(7)). Furthermore, Figure 1(d) and its counterparts illustrate the LDA ratio, emphasizing the equal contributions of covariance and Hessian eigenanalyses to class separability. Importantly, the results also imply that the highest LDA ratio is observed for the combination of the first Hessian eigenvector with the first covariance eigenvector. This observation underscores the significance of projecting data onto the combined space of these primary eigenvectors (as outlined in Eqs (4, 5)), forming the core of our proposed method. These consistent empirical results across multiple datasets not only validate our theoretical premises but also endorse the effectiveness of our proposed method in optimizing class separability.
The evaluation results in Figure 2 compare the performance of different data projection methods, including PCA, KPCA, Hessian, UMAP, LLE, LOL, LDA, KDA, and the proposed method, using 5- or 10-fold cross-validation. It is essential to note that we introduced nonlinearity to the comparative analysis by utilizing distinct kernels determined through grid search in KPCA and KDA for each dataset. Notably, our proposed method consistently outperforms all others, securing the highest average scores across all datasets and evaluation metrics. This consistent superiority of our approach implies its potential as a valuable tool for improving classification performance in various domains, further highlighting its promise as a robust and effective method for dimensionality reduction and data projection.
@@
6 Illustrative case study: WBCD dataset
In this section, we illustrate the practical application of the proposed method using the WBCD dataset. The WBCD dataset comprises 30 predictor attributes for diagnosing breast cancer. This section presents detailed analyses and visualization of the WBCD dataset using our method and other established data projection techniques.
We first examine the separability of classes in the WBCD dataset using basic pairwise attribute projections. Figure 3 shows a pairplot of the first four predictor attributes, colored by class labels. This pairplot projects the 30-dimensional data into 1D and 2D spaces between individual attributes and attribute pairs. As can be visually observed from the figure, these simple projections do not provide good separability between the two classes. This is because these projections are not optimized for class separability. In contrast, established projection methods such as PCA, LDA, and UMAP, as well as our proposed method, are specifically designed and optimized to enhance class separability.
Figure 3: Pairplot of the first four predictor attributes in the WBCD dataset, colored by class labels. This pairplot presents only four of the 30 predictor attributes, projecting the 30-dimensional data into 1D and 2D spaces between individual attributes and attribute pairs. The visualizations show that these projections do not provide good separability between the two classes.
Our proposed method leverages the strengths of both the covariance matrix and the Hessian matrix to achieve optimal class separability. Figure 4 presents the heatmaps of the covariance and Hessian matrices computed from the WBCD dataset. The covariance matrix, computed unsupervisedly from the training set using Eq. 1, is a square and symmetric matrix that captures the statistical spread and relationships between predictor attributes. The Hessian matrix, evaluated on a deep learning model trained on the same training set using Eq. 3, is another square and symmetric matrix that reflects the sensitivity and curvature of the loss function with respect to the model parameters. The DNN test scores are as follows: accuracy: 0.9883, F1 score: 0.9841, AUC ROC: 0.9917, Cohen Kappa score: 0.9749, geometric mean score: 0.9874.
Figure 4: Heatmaps of the covariance matrix and the Hessian matrix for the WBCD dataset. The covariance matrix is computed unsupervisedly from the WBCD training set using Eq. 1. The Hessian matrix is evaluated on a deep learning model trained on the same training set using Eq. 3.
By performing eigenanalysis on the covariance and Hessian matrices, we obtain leading eigenvectors that represent the principal directions of variance and the sharpest curvature, respectively. These eigendirections are crucial for projecting the high-dimensional data into a 2D space that enhances class separability by maximizing the between-class mean distance and minimizing the within-class variances, as described in Theorem 1 and Theorem 2. We project the data into the combined space of the two eigendirections using Eq. 4 and Eq. 5. Figure 5 visually demonstrates the effectiveness of the proposed method by showing the 2D projections of the WBCD dataset. The training set projection highlights the observed visual separation achieved on the training data, while the test set projection confirms good visual separability for new unseen data, underscoring the method’s potential to generalize well. The visual separability observed here will be quantified later using SVM (Figures 7 and 8).
Figure 5: Training set and test set projections on the WBCD dataset using the proposed method. The proposed method projects the high-dimensional data into the leading eigendirections of the covariance and Hessian matrices. The training set projection illustrates how the training data is separated using the proposed method, while the test set projection shows the separation of new unseen data. Visually, the projections indicate good separability between the two classes, demonstrating the effectiveness of the proposed method.
Our proposed method not only enhances class separability but also facilitates interpretability and explainability of the model. Figure 6 presents the parameter contributions to the leading eigenvectors of the covariance and Hessian matrices for the WBCD dataset. The contributions are calculated as the absolute values of the elements in the first eigenvector of each matrix. Parameters are sorted by their absolute contributions, and the horizontal bar plots display these sorted contributions for each parameter. The first plot indicates which attributes contribute the most to maximizing the between-class mean distance, providing insights into the key factors that drive class separation. The second plot shows which attributes contribute the most to minimizing the within-class variances, highlighting the factors that help maintain compact and distinct clusters. These visualizations enable a deeper understanding of the model’s decision-making process, making it easier to interpret and explain the results.
Figure 6: Parameter contributions to the leading eigenvectors of the covariance and Hessian matrices for the WBCD dataset. The contributions are calculated as the absolute values of the elements in the first eigenvector of the covariance and Hessian matrices, respectively. Parameters are sorted by their absolute contributions, and the horizontal bar plots display these sorted contributions for each parameter. The first plot indicates which attributes contribute the most to maximizing the between-class mean distance, while the second plot indicates which attributes contribute the most to minimizing the within-class variances.
Figure 7: Training classification visualizations using various dimensionality reduction methods and linear SVM on the WBCD dataset. The plots show the separation of training data points using PCA, the Hessian method, LDA, UMAP, and the proposed method. Each subplot includes the separating SVM hyperplane and the corresponding F-score, indicating the classification performance of each method. UMAP achieves the highest performance with perfect separation, while the proposed method shows strong performance, surpassing the other methods.
Figure 8: Testing classification visualizations using various dimensionality reduction methods and SVM on the WBCD dataset. The SVM hyperplanes trained on the training set projected using PCA, the Hessian method, LDA, UMAP, and the proposed method, as shown in Figure 7, are then applied to new unseen data points. The plots show the separation of these new data points. Each subplot includes the separating SVM hyperplane and the corresponding F-score, indicating the classification performance of each method on the test set. On new data, the proposed method outperforms all other methods, including UMAP, which had achieved the best performance on the training set. This demonstrates the superior generalization capability of the proposed method.
While Figures 5 illustrate the visual separations achieved by the proposed method, Figures 7 and 8 quantify their degrees of separation. These figures present the classification performance using various dimensionality reduction methods, including PCA, the Hessian method, LDA, UMAP, and the proposed method, evaluated with linear SVM on the WBCD dataset. Each subplot shows the separating SVM hyperplane along with the corresponding F-score, providing a quantitative measure of the classification performance. Figure 7 highlights the separation achieved on the training data, where UMAP achieves the highest performance with perfect separation, while the proposed method demonstrates strong performance, surpassing other methods. Figure 8 shows the performance on new unseen test data, where the proposed method outperforms all other methods, including UMAP, demonstrating its superior generalization capability.